Given that QD= 30 – 5P and QS = 2P -5; also if sales tax of N2 is imposed per unit
i.Calculate the share of tax for both seller and buyer.
What does your answer in (i) suggest in term of elasticity.
1
Expert's answer
2020-09-24T11:05:04-0400

In equilibrium Qd = Qs, so:

"30 - 5P = 2P - 5,"

P = 5,

Q = 5 units.

If sales tax of N2 is imposed per unit, then:

"30 - 5P = 2(P - 2) - 5,"

P = 39/7 = 5.57.

The tax incidence depends on the relative price elasticity of supply and demand. When supply is more elastic than demand, buyers bear most of the tax burden.